OpenStudy (anonymous):

(cot^2u- 1+ csc^2 u) / secu

OpenStudy (anonymous):

What is the question? What is the goal?

OpenStudy (anonymous):

To simplify but using fundemental identities

OpenStudy (anonymous):

By*

OpenStudy (anonymous):

Okay, well then I need to be sure of what it is. \[ \Large \frac{\cot^2(u) - 1 + \csc^2(u)}{\sec(u)} \]

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

First identity to remember is: \[ \large \cot^2(u) + 1 = \csc^2(u) \]

OpenStudy (anonymous):

Okay well i came jp with cotu+cscu+csc^2u but idk if thats right

OpenStudy (anonymous):

This is the pythagorean theorem... comes from \[ \sin^2(u) + \cos^2(u) = 1 \] and dividing the whole thing by \(\sin^2(u)\).

OpenStudy (anonymous):

So is my answer right?

OpenStudy (anonymous):

You want to use that identity to get rid of the \(csc^2(u)\). Can you do that?

OpenStudy (anonymous):

Well csc^2 is also 1/sin^2

OpenStudy (anonymous):

No, what I mean is, replace the csc^2 with cot^2 + 1

OpenStudy (anonymous):

Ohhh ok i see where i messsed up

OpenStudy (anonymous):

Did you get the answer?

OpenStudy (anonymous):

I got (sec)(secu+sin^2+1)???

OpenStudy (anonymous):

Wait no..

OpenStudy (anonymous):

(Cosu)(cosu+cos^2u+1)

OpenStudy (anonymous):

\[ \Large \begin{array}{l} \frac{\cot^2(u) - 1 + \csc^2(u)}{\sec(u)} \\\\ \frac{\cot^2(u)-1+\cot^2(u)+1}{\sec(u)} \\\\ \frac{2\cot^2(u)}{\sec(u)} \\\\ \frac{2\cot^2(u)\cos(u)}{\sec(u)\cos(u)} \\\\ 2\cot^2(u)\cos(u)\\\\ \frac{2\cos^3(u)}{\sin^2(u)} \end{array} \]

OpenStudy (anonymous):

Thank you !!!!!!